OpenStudy (anonymous):

Two thirds of a number added to 6 is equal to the number. What is this number?

OpenStudy (anonymous):

Let's start by: Let x = the unknown number

OpenStudy (anonymous):

0k

OpenStudy (anonymous):

Here's what we're looking at: 2/3 of a number added to 6 is the number

OpenStudy (anonymous):

Let's translate that directly into math.

OpenStudy (anonymous):

Do you remember what the word "of" means?

OpenStudy (anonymous):

Divide

OpenStudy (anonymous):

Not divide... the other one. :)

OpenStudy (anonymous):

(on a side note, usually the word 'per' means divide)

OpenStudy (anonymous):

multiply?

OpenStudy (anonymous):

That's right - multiply! :)

OpenStudy (anonymous):

2/3 of a number added to 6 is the number So we can change the line above to: (2/3) (a number) added to 6 is the number

OpenStudy (anonymous):

Now, what did we decide we were going to call our unknown number?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

That's right. Now, do you remember what the word "is" means?

OpenStudy (anonymous):

"is" is the conjugation of the English verb "to be" for he/she/it or a noun/proper noun.

OpenStudy (anonymous):

So far, we have: (2/3) (x) added to 6 is (x)

OpenStudy (anonymous):

Yes, but what does it mean in the language of mathematics?

OpenStudy (anonymous):

I forgot :(

OpenStudy (anonymous):

What it's equal to?

OpenStudy (anonymous):

Exactly! Is means "="

OpenStudy (anonymous):

So far, we have: (2/3) (x) added to 6 = (x)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now, there's a little trick with the language here. If we have 2/3x added to 6, it would be translated like this: 6 + (2/3)(x) = (x) Remember, if it is "added to" 6 that means we have the 6 first, and we added the 2/3x to it.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

I believe so, yes.

OpenStudy (anonymous):

So, our equation is: 6 + 2/3 x = x

OpenStudy (anonymous):

mm hmm

OpenStudy (anonymous):

So, what would we do to solve this? \[6 + \frac{2}{3} x = x\]

OpenStudy (anonymous):

uhm

OpenStudy (anonymous):

Take away six and have 2/x = x + 6

OpenStudy (anonymous):

Well, if we take away six, we would have: \[\frac{2}{3} x = x-6\]

OpenStudy (anonymous):

Although - instead of that... \[6 + \frac{2}{3} x = x\]Recognize that we have a fraction here.

OpenStudy (anonymous):

Let's clear the fraction by multiplying each piece by 3.

OpenStudy (anonymous):

\[(3)6 + (3)\frac{2}{3} x =(3) x\]

OpenStudy (anonymous):

From there: \[18 + 2x = 3x\]

OpenStudy (anonymous):

18 + 2 + 3z

OpenStudy (anonymous):

Well, the variable would still be with the 2.

OpenStudy (anonymous):

18+2z=3z

OpenStudy (anonymous):

Yes, now what?

OpenStudy (anonymous):

Cover one of 'em up! And you realize that x equals 6! :D

OpenStudy (anonymous):

Not quite...

OpenStudy (anonymous):

18+2z=3z -2z -2z 18=z

OpenStudy (anonymous):

Oh thanks

OpenStudy (anonymous):

do you believe in god?

OpenStudy (anonymous):

So the number we're looking for is 18. :) You're welcome!!

OpenStudy (anonymous):

Do you believe in God?

OpenStudy (anonymous):

Does my medal depend on my answer to that question?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

My answer is 'no'